OpenStudy (anonymous):

A man is walking down the beach away from lighthouse A and toward point C. When he is x kilometers away from lighthouse A and has not yet reached point C, write the illumination he receives as a function of x in simplest form.

OpenStudy (anonymous):

Lighthouse A, located on an ocean shore, uses a 10-kW light. Lighthouse B uses a 20-kW light and is located 8 km out to sea from a point 6 km down the beach from Lighthouse A.

OpenStudy (anonymous):

you can measure the distance from point B using pythagorean thm d=82+(6−x)2−−−−−−−−−−−√=x2−12x+100−−−−−−−−−−−−√ is illumination measured by watts*distance ? if so then function is l=10x+20x2−12x+100−−−−−−−−−−−−√

OpenStudy (anonymous):

:) have to go - did this help?

OpenStudy (anonymous):

Thanks. The formula is 10/x^2+20/sqrt(x^2-12x+100)

OpenStudy (anonymous):

what should i do from now?

OpenStudy (anonymous):

was hoping I could paste the doc in Word to you - shoot I have to go - :( I am sorry I couldn't help you more )

OpenStudy (anonymous):

let me see if I can try again

OpenStudy (anonymous):

I can forward this to a friend of mine in FB later on if noone helps you (he is a math teacher in California) and he can respond - his FB is: Just Plane Math on Facebook.

OpenStudy (anonymous):

lol I am a girl ;)

OpenStudy (anonymous):

https://www.facebook.com/pages/Just-Plane-Math/255647914493669 That is his page - and he is usually able to help when others can't :)

OpenStudy (anonymous):

:) thanks

OpenStudy (anonymous):

K, he is responding - I will try to copy this to you-